Summit Systems has an equity cost of capital of 11.5 %​, will pay a dividend of ​$1.75 in one​ year, and its dividends had been expected to grow by 5.5 % per year. You read in the paper that Summit Systems has revised its growth prospects and now expects its dividends to grow at a rate of 3.0 % per year forever. a. What is the drop in value of a share of Summit Systems stock based on this​ information? b. If you tried to sell your Summit Systems stock after reading this​ news, what price would you be likely to​ get? Why? a. What is the drop in value of a share of Summit Systems stock based on this​ information? The drop in value of a share of Summit Systems stock is ​$ nothing. ​(Round to the nearest​ cent.)
Answer #1

Original share price = D1/(k-g)

=1.75/(11%-5.5%)

=31.8182

Price after news = 1.75/(11%-3%)=21.875

Hence Drop in share price = 31.8182-21.875

= $9.94

B: if you tried to sell the stock off after this news is out, you will likely get the reduced price because the market expectations have changed and the share price will be determined according to the new growth rate.